OpenStudy (anonymous):

. A company builds computers. It costs $6,700 to build 10 computers and $12,200 to build 20 computers. Which equation models the cost, C(x), as a linear function of the number of computers built, x? A. C(x) = 550x -1200 B. C(x) = 550x + 1200 C. C(x) = 1200x - 550 D. C(x) = 1200x + 550

OpenStudy (anonymous):

i will tell you the slope is 12200-6700 divided by 20-10 5500/10 m = 550 use the formula y - y1 = m (x - x1) to find your line x1 = 10 y1 = 6700

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

which is the answer? It is B right?